Problem # 2: Suppose that Y , Y,,...,Y, denote a random sample of size n from a normal population with mean u and variance oSuppose that Y1 , Y2 ,..., Yn denote a random sample of size n from a normal population with mean μ and variance  2 .
Answer #1

Be a ay Jet xn Ayamma (ws, Ma> : pdf f*: $4(e)= } ***, zro Map of X: MyC) = E(eta). Jak si zat de whee, te=484): para cena ŞeWe will use the info above

G Note: Y= (1-28 n te ne za Sauna ( 1 ,2) - E(Y)- MI42=(1-1) a) E (8-1952 ) -(mv) CD 1) E(62) = (2-1) [: nul, cine Coretant]